The sum of three consecutive integers is 24. What is the product of these three integers?
Answer by stanbon(75887)

The sum of three consecutive integers is 24. What is the product of these three integers?
-----
1st: x-1
2nd: x
3rd: x+1
----
sum = 24
3x = 24
x = 8
---
1st: x-1 = 7
2nd: 8
3rd: 9
----
Product = 7*8*9 = 504
==============================
Cheers,
Stan H.
=====================